Elemental calcium is produced by the electrolysis of molten CaCl2.

A) What mass of calcium can be produced by this process if a current of 6600 A is applied for 50 h? Assume that the electrolytic cell is 66 % efficient.

Express your answer using two significant figures.

*incorrect answer = 8.1 x 10^8 g; 8.6 x 10^8 g

B) What is the minimum voltage needed to cause the electrolysis?

Express your answer using three significant figures.
Answer #1

A)

w = Act / ZF

A = 40.078 g/mol; c = 6600 A; t = 50 h = 180000 s; Z = +2 e-; F = 96485 C/mol of electrons

w = [40.078 g/mol x 6600 A x 180000 s] / [2 e- x 96485 C/mol of electrons] = 246736.093693 g

but cell is 66% efficient = 246736.093693 x (66 / 100) = 162845.821838 g = 1.6 x 105 g

Answer: 1.6 x 105 g

B)

Ca2+(aq) + 2e → Ca(s) E0red = -2.87 V,

Cl2(g) + 2e → 2Cl(aq) E0red = +1.36 V,

E0cell = 1.36 + 2.87 V = 4.23 V

Answer: the minimum voltage needed to cause the electrolysis = 4.23 V

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")